Ammonium chloride crystallizes in a bcc lattice with edge length 387 pm . If the size of cl is 181 pm size of NH4+ ion would be ?
116
154
206
174

Grade:12

1 Answers

Vasantha Kumari
9 years ago

Here, NH4Cl has BCC cell structure and ions touch each other along the body diagonal of the cube.

Thus, 2r+ + 2r =v3 a that is

2(r+ + 181 pm )= =v3 (387 pm)

(r+ = v3/2 (387 pm)–181 pm = 154.15 pm

The size of NH4+ ion is 154.15 pm

Thanks & Regards,

Vasantha Sivaraj,

askIITians faculty